OpenStudy (mrmanman305):

Let u = <-5, -9>, v = <6, 8>. Find -8u - 2v. HELP!!!

OpenStudy (bobo-i-bo):

-8u-2v=-8<-5,-9>-2<6,8>

OpenStudy (bobo-i-bo):

A property of vectors: a<b,c>=<ab,ac>

OpenStudy (bobo-i-bo):

From the stuff I have given you, can you solve it now? Or do you not understand what i'm on about?

OpenStudy (mrmanman305):

So its kinda like distributive property

OpenStudy (mrmanman305):

I don't get how to solve it lol

OpenStudy (freckles):

a<b,c>=<ab,ac> means the components of <b,c> are each going to be multiplied by a ab means a times b ac means a times c

OpenStudy (bobo-i-bo):

For example, 5<4,6>=<4*5,6*5>=<20,30>

OpenStudy (bobo-i-bo):

Also, another property of vectors you need is this: <a,b>+<c,d>=<a+c,b+d> So for example: <1,2>+<3,4>=<1+3,2+4>=<4,6>

OpenStudy (mrmanman305):

thank you my good people